OpenStudy (anonymous):

PLZ help gold medal will b awarded.write the equation of the line that is parallel to the line y=2x+2 and passes through the point (5,3)

OpenStudy (austinl):

If a line is parallel that means that it shares the same slope. In our case, 2. Now all we need to do is find the y-intercept which we can do by plugging in the values of x and y given by the coordinate. Do you think you can solve from here?

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

hey can i get answer for this one cuz i got to go eat

OpenStudy (anonymous):

plz

OpenStudy (austinl):

We have the given equation, \(y=2x+2\) We know that our line will be parallel, and therefore have the same slope. \(y=2x+b\) Now we need to find b, we do this by inputting the values for x and y from the coordinate given. \((x,y)\rightarrow (5,3)\) \(3=2(5)+b\) \(3=10+b\) \(b=-7\) Therefore we have, \(y=2x-7\) Does this make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thx